a retangler garden has a length that is nine less than three times its width. if the perimeter is 89 feet, what are the dimensions.
Found 2 solutions by 2897696, stanbon:
Answer by 2897696(96)

width: x
length:3x-9
perimeter: width +width +length+ length=89
or ------ 2width+2length=89
2x+2(3x-9)=89
2x+6x-18=89
8x=107
x=13.375
width :13.375 ft
length:
3(13.375)-9
31.125 ft
ANSWER length:31.125ft
width 13.375 ft

Answer by stanbon(75887)

a retanguler garden has a length that is nine less than three times its width. if the perimeter is 89 feet, what are the dimensions.
------------------------------
Let width be "x".
Then length = 3x-9
------
Perimeter = 2(length + width)
89 = 2(3x-9+x)
89 = 2(4x-18)
44.5 = 4x-18
4x = 62.5
x = 15.625 ft. (width)
3x-9 = 37.875 (length)
============================
Cheers,
Stan H.